OpenStudy (anonymous):

What is the general form of the equation of a circle with its center at (-2,1) and passing through (-4,1)? x^2 + y^2+4x-2y+1=0 x^2 + y^2+4x-2y+9=0 x^2 - y^2+2x+y+1=0

OpenStudy (anonymous):

\[ (x+2)^2+(y-1)^2=r^2 \] We still have to fine \( r^2\).

OpenStudy (anonymous):

The distance from the center to a point on the circle is r and hence r=2

OpenStudy (anonymous):

\[ (x+2)^2+(y-1)^2=4\\ 4 + 4 x + x^2 + y^2-2 y+1=4\\ x^2+4 x+y^2-2 y+5=4\\ x^2+4 x+y^2-2 y+1=0 \]

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

so just saw your new post

OpenStudy (anonymous):

It is the first choice

OpenStudy (anonymous):

Did you understand it?

OpenStudy (anonymous):

i just fiquired out it is the 1st choice. x^2 + y^2+4x-2y+1=0 this might kill me. can you check my answer on another one?

OpenStudy (anonymous):

Post is as new problem and tag me

OpenStudy (anonymous):

ok thanks i can't open it up foer some reasom issues with open study

OpenStudy (anonymous):

You have to close this one and open a new one

OpenStudy (anonymous):

See the picture